OpenStudy (anonymous):

mrs. boyd has a desk full of quarters and nickels. if she has a total of 26 coins with a total face value of $5.10, how many of the coins are nickels?

OpenStudy (slaaibak):

How much is a nickel?

OpenStudy (turingtest):

0.25x+0.05y=5.1 x+y=26 solve the system

OpenStudy (slaaibak):

There are x= quarters and 26 - x nickels. so 0.25x + (26 - x) * [VALUE OF NICKEL] = 5.1

OpenStudy (anonymous):

7 nickels

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

ty everyone